Problem 2.

Compute the dielectric constants of the media in the following scenarios:

i. The electric potential due to a 1∗10−9C charge 1 meter away on a test charge is 9 volts. What might be the medium in this interaction?

ii. The electrostatic energy is 2.93∗10−20J for two electrons one angstrom apart. What might be the medium in this interaction?
Answer #1

1)

Electric Potential, V is given by

Ο 1 V Απερε, Τ κ ----(1)

Where Q is the charge, r is the distance

The value of

1 Απερ-

is 9*109

Rearranging the equation (1) we get,

E,=9 109 Vr

Substituting the above values we get,

\varepsilonr = 9*109*1*10-9/(9*1)

=> \varepsilonr = 1

This is the relative permittivity of vacuum. Therefore, the medium is vacuum (also called as free space)

ii) Electrostatic energy, U(r) is given by

Τ υr)- κ Απερε, Υ ------(2)

Rearranging equation (2) we get,

E9 109 rU (r)

Substituting the value of q = Q = 1.6x10-19 C (which is the value of charge of electron), r = 1x10-10 m, U=2.93x10-20 J we get,

\varepsilonr = 9*109*1.6*10-19*1.6*10-19/(1*10-10*2.93*10-20)

\varepsilonr = 78.6

This value is the relative permittivity of pure water. Therefore the medium in this case is pure water.